# Using Knowledge Graphs to harvest datasets for efficient CLIP model training

### v1

@@ 1 Introduction
Figure 1: We demonstrate how to harvest datasets for training CLIP models with an improved quality-cost trade-off, either for a generic domain (top) or an expert domain (bottom).
@@
We built a dataset we named EntityNet, where we leveraged knowledge graphs and targeted web image search. Specifically, from the knowledge graphs Wikidata and WordNet, we collected 135k entities (e.g. eagle) as well as their aliases and descriptions. We extracted entity attributes from Wikidata related to color, partonomy, behavior, and other aspects, which we then used to guide an LLM in generating entity-attribute queries for image search.. For example, from the entity plastic and the attribute small we generated the search term small plastic item.
@@
Training on this dataset, we obtain a foundation model that is both specialized on the target expert domain and is also able to understand the overall visual world. In our domain-specific evaluations on iNaturalist and RareSpecies, the model demonstrates robust generalization and clearly surpasses CLIP models trained on much more data (Figure 1). On ImageNet, we demonstrate our dataset to be highly compute efficient and to achieve a performance comparable to models trained 20x longer (Figure 1).
@@ 2 Related work
Datasets. Many recent works investigated ways to build large-scale datasets for multimodal training. Radford et al. [29] train the original CLIP model on a closed set of 400M images, with the model weights being released, but not the data. They build their dataset by collecting image-text pairs, where the text includes frequent terms derived from Wikipedia or WordNet nouns, and approximately balance the classes in the result. As a first approach to create a public dataset of this size, Schuhmann et al. [33] built a dataset with 400 million image-text pairs by filtering HTML data from Common Crawl (LAION-400M) [30]. Their main method of filtering is to remove all image-text pairs that have less than 0.3 similarity estimated by the CLIP model. In a follow-up work, they [34] scaled their approach up one order of magnitude with the multilingual LAION-5B dataset. Xu et al. [48] intended to replicate the original CLIP’s data curation approach. They collected image-text pairs from CommonCrawl and filtered them using Wikipedia and WordNet, then balanced the results. Gadre et al. [10] proposed DataComp, a filtering challenge with a candidate pool of up to 13B image-text pairs from CommonCrawl, where the goal is to filter this candidate pool and to run a fixed training pipeline on the resulting data. They proposed a baseline DataComp-1B dataset with 1.4B pairs filtered with a combination of CLIP score and clustering CLIP embeddings to find images close to ImageNet [7] training examples. Fang et al. [8] trained a Data Filtering Network on an internal dataset of 357M human-verified image-text pairs and finetuned it on a set of public, human-annotated datasets. They filtered 42B candidates into the DFN-5B dataset and trained the current top model of the OpenCLIP leaderboard [14].
These large datasets have mostly replaced smaller datasets like ConceptualCaptions12M (CC12M) [4], which relies on unimodal heuristics and Google Cloud Vision APIs to predict image-text similarity. Another popular small dataset is Yahoo Flickr Creative Commons 15M (YFCC15M), a subset of 15M image-text pairs obtained from the YFCC100M dataset which is based on Flickr [38]. Although many works are mostly concerned with scaling up multimodal datasets and models as much as possible, we aim to improve research on high-quality CLIP models in scenarios where data and compute efficiency is important, for example, for setting up a CLIP model for an expert domain or for scientific analysis of CLIP training.
Stevens et al. [35] aimed for a general vision model named BioCLIP for organismal biology and curated the TreeOfLife-10M dataset based on the Encyclopedia of Life [1], iNaturalist 2021 [41] and BIOSCAN-1M [11]. BioCLIP is trained on a mix of english and latin entity names. For evaluation, they curated the RareSpecies benchmark, which tests generalization to 400 species unseen during training. While the authors of BioCLIP used biological domain knowledge to build their dataset, we rather rely on knowledge graphs and propose a dataset collection method for arbitrary domains.
@@
Training algorithms. Various works are concerned with improving CLIP from the algorithmic side. Li et al. [16] simply train on low resolution first, then finetune on higher resolution later. In another work [17], the authors additionally mask a substantial portion of the image to further reduce the amount of input during training. Zhai et al. [51] propose using a sigmoid loss which reduces the computational load especially in big distributed settings. They follow up [39] by extending the training objective using multiple previously developed techniques, including captioning-based pretraining [45], self-distillation [24] and online data curation [40] into one unified recipe. Vasu et al. [42] enhance their training data with synthetic captions created by an image captioning model and use an ensemble of CLIP teachers to train their model. This way, they can increase the learning efficiency by transferring knowledge from bigger models to their smaller models. Chen et al. [5] evaluate the choice of vision encoder and design a hybrid architecture that improves over vanilla vision transformers when trained under the CLIP framework. Such algorithmic improvements are orthogonal to our research, since they would potentially also improve training on our dataset. In this work, we focus on data improvements and fix the algorithm and architecture choices, since this also allows to easily and fairly compare to a big set of already trained Vision Transformer (ViT) based CLIP models.
@@
Li et al. [18] scale down CLIP and analyse the influence of different data, architecture, and training strategies. They find that especially large models need larger datasets, and data quality plays an important role. They create higher quality datasets by applying CLIP filtering to the 3.4B WebLI dataset [6], while we aim to use a different dataset collection process.
@@ 3 Dataset creation
In this section, we outline our dataset creation process, consisting of four main steps: entity extraction, attribute generation, query building, and image search. This process is generally applicable to all visual domains covered by the underlying knowledge graph. In this work, we construct a dataset covering the majority of common visual entities in our world. Within this broad domain, we particularly focus on animals and plants, referred to as the living subset in this section. See Figure 2 for an illustration of the dataset creation process.
@@ 3.1 Entity extraction
A high-quality list of visual entities forms the basis for our dataset. We build our list of entities from the Wikidata knowledge graph [43], taking advantage of the hierarchical structure provided by the subclass of relation within Wikidata. For example, the entity dog is a subclass of the pet entity, which in turn is a subclass of domesticated animal entity. Through this hierarchy, we can easily collect all entities under a particular governing or super-entity. First, we manually build a list of 21 super-entities that cover most physical and visual entities in Wikidata. For the living subset, the super-entities are just animal and plant. Examples of non-living super-entities are food, building, or physical tool, with all super-entities listed in the supplementary material. Next, we extract all entities from Wikidata that are linked to at least one of the super-entities through the subclass of relation. For animals and plants, Wikidata also models their biological taxon hierarchy via the parent taxon relation. Because the taxon hierarchy substantially increases the coverage of our living subset, we use it together with the regular subclass hierarchy to extract living entities. Note that we do not want to include named entities (e.g., specific persons). Since Wikidata models these via the instance of relation, our extraction process excludes them by using only the subclass of and parent taxon relations. For every entity, we also download its name, description, aliases, and its number of Wikimedia sitelinks  [^1] The number of Wikimedia sitelinks is a commonly used and high-quality proxy for the popularity of an entity [26]. as additional information. See Table 1 for examples of entities. Finally, we apply two filtering steps: First, we remove all entities with a sitelink count below a predefined threshold, eliminating very rare or low-quality entities which are unlikely to produce strong search results. We find that this filtered list of entities still contains a significant amount of non-visual entities. To address this, we use a LLM to classify each entity as visual or non-visual, discarding the non-visual ones.
For our expert domain, the living subset, we also add all nouns from WordNet [9] that are a subclass of the living thing node, excluding humans, named entities and entities that cannot be seen with the bare eye, e.g., microorganisms.
Finally, we employ heuristic methods to detect and remove potentially offensive entities via a profanity filter.
Table 1: Examples of entities and additional information as extracted from the Wikidata knowledge graph. Name, description, and aliases are used as text labels during training. The number of sitelinks are considered a proxy for an entity’s popularity. The name and aliases are used during search.
@@
3.2 Attribute generation
Besides searching for the entities directly, we also aim to search for variations of them in different contexts, by combining them with attributes. We manually define 6 visual attribute categories we want to search for: Color, Pattern and texture, Parts, Shape and size, Environment, and Other. We extract potential attributes for each entity from the Wikidata knowledge graph and prompt an LLM  [^2] In particular, we use three LLMs and merge their generated attributes: Qwen2.5 7B [49], OpenAI GPT-4o, and OpenAI GPT-4o mini (both accessed via the API at platform.openai.com) with this entity and attribute information to generate a list of visual attributes. For each attribute we also generate a search query combining the attribute itself with the corresponding entity. See Table 2 for examples of generated attributes and search queries. We generate between 1 and 10 attributes per category and generate them for the most popular entities only, as image search engines fail to respect attributes in search queries for rare entities, where they often even struggle to return good results for the entity alone.
Table 2: Examples of attributes and corresponding search queries for different entities as generated by the LLM. Search queries are used for image search and during training.
@@ 3.3 Query building
For the entities themselves, we use their names and aliases as search queries. We search entity-attribute combinations using the search queries generated by the LLM. We then build another set of queries based on the attributes: First, we determine the natural type for each entity. The natural type of an entity is the super-entity that a human would most likely associate with it, e.g., bird for eagle, or clothing for hat. It is neither too general nor too specific, and can typically be used to disambiguate the entity from other entities with the same name. Here, we use an LLM to select the most fitting super-entity from an entity’s super-class hierarchy as its natural type. We also let the LLM generate a short description as to why the selected natural type is a good fit for the entity and use it during training as a potential text label. We then replace entity mentions in the attribute search queries with their natural types. For example, the attribute query eagle in its nest may turn into bird in its nest, or black BMW M4 into black car.
@@ 3.4 Image search and filtering
We execute our search queries using the image search APIs of Bing and Google. Initial tests on the living subset revealed Bing’s search results to be of much higher quality at a lower cost, so we rely solely on the Bing API for all other queries. The image search APIs also provide the URL for the website hosting the image, which we use to collect alt texts from the HTML image tag. After downloading images and alt texts, we perform the following postprocessing steps:
  • Similar to Changpinyo et al. [4], we apply relaxed filtering heuristics. We do not use any multimodal filtering but instead rely on the search engines to provide image-text correspondences. We remove text that is longer than 500 chars or formatted in JSON. We also remove images with an aspect ratio of more than 4 or covering less than 4096 pixels.
  • We deduplicate all downloaded images using the Self-Supervised Descriptor for Image Copy Detection method (SSCD) [27]. For duplicates, we retain the largest image and collect all unique alt texts from the duplicates.
  • We detect duplicates between the images and all evaluation datasets using the same SSCD method.
Our final dataset comprises approximately 33M images and 45M alt texts, obtained from 416k queries. This amounts to 79 images per query and 1.4 alt texts per image on average. The total cost for all image search API calls was around 10,000\$. See Table 3 for an overview over the building blocks and proportions of our dataset.
@@
World entity + attribute 19M 139k 6k 16k 20k 16M small handbag
@@
Living entity + attribute 9M 53k 1k 3k 5k 6M tropical plant
@@ 4.1 Training
We trained all models using the standard CLIP loss [29] with a batch size of 8,192 for pretraining and 32,768 during finetuning, along with random resized crop augmentation. We sampled text labels from both the image alt texts and knowledge graph - for each image, 50% of the time we chose a random alt text for a given image, and 50% of the time we chose randomly between search query, aliases, or descriptions of the corresponding entity. Our training code is based on OpenCLIP [14]. Further training details and an example for the text sampling can be found in the supplementary material. We trained all models for 18 epochs using AdamW [23]. Training on 33M images takes $\sim$55 hours on 8 L40s GPUs with 48GB VRAM per GPU.
@@ 4.2 Evaluated models
On our EntityNet dataset, we trained ViT CLIP models of size B-32 and B-16 from random initialization. For a comparison with a similarly sized dataset, we also trained models on CC12M by downloading all available URLs, and then detecting and removing duplicates relative to the evaluation datasets using the same procedure as detailed in Section 3.4, obtaining 9.3M images. We finetuned B-32 and B-16 CLIP models trained on DataComp-1B on our dataset to compare finetuning and pretraining performance. In addition to the models trained on EntityNet and CC12M, we evaluate the original OpenAI CLIP [29] and models pretrained on DataComp-M/L/1B, CommonPool-M/L [10], and DFN-5B [8]. We also compare with the domain expert model BioCLIP [35], a ViT-B-16 CLIP model finetuned from OpenAI-CLIP on 10M biological image-text pairs.
@@ 4.3 Object classification evaluation
To test the VLMs on object classification we use the same procedure as CLIP [29]. Given an image $I$, class names $C_{1},...,C_{N}$, image encoder $f$ and text encoder $g$, we embed the image using the image encoder $\mathbf{v}=f(I)$. To acquire a text embedding for class $C_{c}$, the CLIP authors started by directly encoding the class names as $\mathbf{w}_{c}=g(C_{c})$, e.g., dog. Alternatively, they created several prompts $P$ using templates, e.g., graffiti of a dog, a photo of the cool dog, etc., then encoded each prompt, and computed the average embedding: $\mathbf{w}_{c}=\sum_{p\in P}g(p)/|P|$. They referred to this approach as using “context prompts”. Finally, given the image and text embeddings, the prediction $p$ is the class which has the highest cosine similarity to the image. We evaluate all models on just encoding the class name, as well as on using the average embedding of the 80 context prompts that the CLIP authors used for ImageNet, and report the higher top-1 accuracy. For zero-shot object classification, we require models to not have trained on the training set of the benchmark, in order to test “generalization to unseen datasets” [29]
@@
Benchmarks in the generic domain. We evaluate on ImageNet [7], a popular image classification benchmark [32]. We use the ILSVRC2012 validation set, which contains 50,000 images from 1,000 classes. The classes include simple objects such as park bench, but also more fine-grained labels like 23 types of terrier dogs, e.g., Staffordshire Bull Terrier. We further evaluate the robustness under distribution shifts on ImageNet-A [13], ImageNet-R [12], ImageNet-Sketch [46], ImageNet-V2 [31], and ObjectNet [3] as proposed by Taori et al. [36]. ImageNet-A contains 7500 samples of 200 ImageNet classes. The samples were adversarially filtered to make ResNet-50s misclassify them, providing a more challenging test. ImageNet-R contains 30000 renditions, such as paintings or embroidery, of 200 ImageNet object classes. ImageNet-Sketch contains 50000 sketches, covering 200 ImageNet classes. ImageNet-V2 replicates the original ImageNet generation process, providing an additional 10000 test images. The object-centered ObjectNet contains 18574 images from 113 ImageNet classes with control over background, rotation and viewpoint.
@@
Benchmarks in the expert domain. We evaluate our models on iNaturalist 2021 [41], a fine-grained species classification benchmark that contains 100k images in the validation set of 10k different species. Similar to Parashar et al. [25], we test models on both the english common name and the latin taxon name. We report the best results over both languages. We further test on the Caltech-UCSD Birds (CUB) [44] dataset, which contains 5,794 images of birds in the original author’s test set, each annotated as one of 200 fine-grained bird species, e.g., grasshopper sparrow. Additionally, we evaluate on the Rare Species benchmark proposed by Stevens et al. [35], that comprises 400 species with 30 images each specifically tailored to assess generalization to unseen taxa. To comply with the benchmark requirements of not seeing the testing 400 species during training, we exclude all entities and queries from our dataset that appear in RareSpecies, using substring matching. As class names, we evaluate all text types proposed by Stevens et al. [35]: combinations of the latin taxonomy and the english common name. Same as in Section 4.3 we evaluate on both the CLIP ImageNet prompt and no prompt, and report the better of both accuracies.

### v3

# Code and data: https://entity-net.github.io

@@ 1 Introduction
Figure 1: We demonstrate how to harvest datasets for training CLIP models with an improved quality-cost trade-off, for a generic (left) or an expert domain (right).
@@
We built a dataset we call EntityNet, where we leveraged knowledge graphs and targeted web image search. Specifically, from the knowledge graphs Wikidata and WordNet, we collected 135k entities (e.g. eagle) as well as their aliases and descriptions. We extracted entity attributes from Wikidata related to color, partonomy, behavior, and other aspects, and used them to guide an LLM in generating entity-attribute queries for image search. For example, from the entity plastic and the attribute small we generated the query small plastic item.
@@
Training on this dataset, we obtain a foundation model that is both specialized on the target expert domain and is also able to understand the overall visual world. In our domain-specific evaluations on iNaturalist and RareSpecies, the model demonstrates robust generalization and clearly surpasses CLIP models trained on much more data (Figure˜1). On ImageNet, we demonstrate our dataset to be highly compute efficient and to achieve a performance comparable to models trained 20x longer (Figure˜1).
@@ 2 Related work
Datasets. Many recent studies have investigated methods for building large-scale datasets for multimodal training. Radford et al. [1] trained the original CLIP model on a private dataset of 400M images using image-text pairs with text derived from Wikipedia and WordNet terms. Schuhmann et al. [10] further built the publicly available LAION-400M dataset by filtering HTML data from Common Crawl  [11] based on the similarity estimated by the CLIP model. In a follow-up work, they [12] scaled their approach one order of magnitude with the multilingual LAION-5B dataset. Xu et al. [13] sought to replicate the original CLIP’s data curation approach. Gadre et al. [6] proposed DataComp, a filtering challenge containing up to 13B image-text pairs from CommonCrawl, and a baseline DataComp-1B dataset. It contains 1.4B pairs filtered with a combination of CLIP score and clustering to match ImageNet [14] training examples. Fang et al. [7] trained a Data Filtering Network on 357M human-verified image-text pairs, which they used to filter 42B candidates into the DFN-5B dataset and use that dataset to train the current top model of the OpenCLIP leaderboard [15]. These large datasets have largely supplanted smaller ones like ConceptualCaptions12M (CC12M) [16], relying on unimodal heuristics, and Yahoo Flickr Creative Commons 15M (YFCC15M), a derived subset of 15M image-text pairs from Flickr [17]. While many prior works have focused on scaling up multimodal datasets and models, we aim to improve research on high-quality CLIP models when data and compute efficiency is essential, such as setting up a CLIP model for an expert domain or for scientific analysis of CLIP training.
Stevens et al. [18] curated the TreeOfLife-10M dataset from biological sources [19, 20, 21] to train BioCLIP, a model for organismal biology. They evaluated it on RareSpecies, a benchmark of 400 species not seen during training. While BioCLIP leverages domain-specific biological knowledge, we propose a dataset construction method that generalizes to arbitrary domains using knowledge graphs.
@@
Training algorithms. Several works have investigated algorithmic improvements to CLIP. Li et al. [22] simply train and fine-tune with different image resolutions, while Li et al. [23] suggests masking parts of the image to reduce computation. Zhai et al. [24] propose a sigmoid loss which reduces the computational load especially in big distributed settings. They follow up [25] by extending the training objective using multiple previously developed techniques, including captioning-based pretraining [26], self-distillation [27] and online data curation [28] into a unified training strategy. Vasu et al. [29] improve learning efficiency with synthetic captions created by an image captioning model and an ensemble of CLIP teachers to train their model. Chen et al. [30] evaluate vision encoder choices and design a hybrid architecture that improves over vanilla vision transformers (ViT) based CLIP models. These algorithmic improvements are orthogonal to our contribution. In this work, we fix the algorithm and architecture to enable a fair comparison with ViT-based CLIP baselines.
@@
Li et al. [8] analyze scaling effects across data, architecture, and training strategies, showing that huge models require larger datasets, and data quality plays a crucial role. They create improved datasets by filtering the 3.4B WebLI dataset [31] with CLIP, while we pursue a different dataset collection process.
@@ 3 Dataset creation
Our dataset creation process consists of four steps: entity extraction, attribute generation, query building, and image search. This process is generally applicable to all visual domains covered by the underlying knowledge graph. We construct a dataset covering most visual entities in our world, additionally focusing on animals and plants, referred to as the organism subset. See Figure˜2 for the dataset creation process and Table˜1 for examples of entities and attributes.
@@ 3.1 Entity extraction
A high-quality list of visual entities forms the basis for our dataset, built from the Wikidata knowledge graph [32] and utilizing the hierarchical structure provided by the subclass of relation within Wikidata. For example, the entity dog is a subclass of the pet entity, which in turn is a subclass of domesticated animal. This hierarchy enables easy collection of entities related to a super-entity. First, we manually build a list of 21 super-entities that cover most physical and visual entities in Wikidata. For the organism subset, the super-entities are just animal and plant. Examples of non-organism super-entities include food, building, or physical tool, with all super-entities listed in the supplementary material. Next, we extract all entities from Wikidata linked to at least one of the super-entities through the subclass of relation. For animals and plants, Wikidata also models their biological taxon hierarchy via the parent taxon relation. Because the taxon hierarchy substantially increases the coverage of our organism subset, we use it together with the regular subclass hierarchy to extract entities. We exclude named entities (e.g., specific persons), as Wikidata models these via the instance of relation; we focus solely on the subclass of and parent taxon relations. For every entity, we also download its name, description, aliases, and its number of Wikimedia sitelinks  [^1] The number of Wikimedia sitelinks is a commonly used and high-quality proxy for the popularity of an entity [33]. as additional information. Finally, we apply two filtering steps: First, we remove all entities with a sitelink count below a predefined threshold, eliminating very rare or low-quality entities unlikely to produce strong search results. We then use a LLM to filter out any remaining non-visual entities.
For our expert domain, the organism subset, we also add all nouns from WordNet [34] that are a subclass of the living thing node, excluding humans, named entities and entities that cannot be seen with the bare eye, e.g., microorganisms. Finally, we employ heuristic methods to detect and remove potentially offensive entities via a profanity filter.
Table 1: Top: Examples of entities and additional information as extracted from the Wikidata knowledge graph. Bottom: Examples of attributes and corresponding search queries for different entities as generated by the LLM.
@@
3.2 Attribute generation
Besides searching for the entities directly, we also aim to search for variations of them in different contexts, by combining them with attributes. We manually define 6 visual attribute categories we want to search for: Color, Pattern and texture, Parts, Shape and size, Environment, and Other. We extract potential attributes for each entity from the Wikidata knowledge graph and prompt an LLM [^2] We use three LLMs and merge their generated attributes: Qwen2.5 7B [35], OpenAI GPT-4o, and OpenAI GPT-4o mini (both accessed via API at platform.openai.com) with this entity and attribute information to generate a list of visual attributes. We first considered generating attributes without categories, however, the results lacked diversity, and adding categories improves the variety of attributes. For each attribute we also generate a search query combining the attribute itself with the corresponding entity. We generate between 1 and 10 attributes per category and generate them for the most popular entities only, as image search engines fail to respect attributes in search queries for rare entities, where they often even struggle to return good results for the entity alone.
@@ 3.3 Query building
For the entities themselves, we use their names and aliases as search queries. We search entity-attribute combinations using the search queries generated by the LLM. We then create additional queries based on the attributes: First, we determine the entity’s natural type – the super-entity a human would most likely associate with it, e.g., bird for eagle, or clothing for hat. It is neither too general nor too specific, and can typically help disambiguate entities sharing the same name. We use an LLM to select the most fitting super-entity from an entity’s super-class hierarchy as its natural type and generate a brief description explaining why this type is appropriate. The description is used during training as a potential text label. We then replace entity mentions in the attribute search queries with their natural types. For example, the attribute query eagle in its nest may turn into bird in its nest, or black BMW M4 into black car.
@@ 3.4 Image search and filtering
We execute our search queries using the image search APIs of Bing and Google. Initial tests on the organism subset revealed Bing’s search results to be of much higher quality at a lower cost, so we rely solely on the Bing API for all other queries. The image search APIs also provide the URL for the website hosting the image, which we use to collect alt texts from the HTML image tag. After downloading images and alt texts, we perform the following postprocessing steps.
Similar to Changpinyo et al. [16], we apply relaxed filtering heuristics. We do not use multimodal filtering, but rely on search engines to provide image-text correspondences. We remove JSON-like and too long text. We also remove images with an aspect ratio of more than 4 or covering less than 4096 pixels.
We deduplicate all downloaded images using the Self-Supervised Descriptor for Image Copy Detection method (SSCD) [36]. For duplicates, we retain the largest image and collect all unique alt texts and related entities from the duplicates. Deduplication increases the dataset diversity per sample, since the domain coverage stays the same, while the number of samples decreases. We also remove images that appear in any evaluation dataset using the same SSCD method.
Our final dataset comprises approximately 33M images and 45M alt texts, obtained from 416k queries. This amounts to 79 images per query and 1.4 alt texts per image on average. The total cost for all image search API calls was around 10,000\$. See Table˜2 for an overview over our dataset.
@@
+ attribute 19M 139k 6k 16k 20k 16M small handbag
@@
+ attribute 9M 53k 1k 3k 5k 6M tropical plant
@@ 4.1 Training
We trained all models with the standard CLIP loss [1], using a batch size of 8,192 for pretraining and 32,768 during finetuning, along with random resized crop augmentation. We sampled text labels from both the image alt texts and the knowledge graph. For each image, 50% of the time, we chose a random alt text, and 50% of the time, we chose randomly between search queries, aliases, or descriptions of the corresponding entity. We trained all models for 18 epochs. Training on 33M images takes $\sim$55 hours on 8 L40s GPUs (48GB VRAM per GPU). Our training code is based on OpenCLIP [15]. Further training details and text sampling examples are in the supplementary material.
@@ 4.2 Evaluated models
On our EntityNet dataset, we trained ViT CLIP models of size B-32 and B-16 from random initialization. For a comparison with a similarly sized dataset, we also trained models on CC12M by downloading all available URLs, and then detecting and removing duplicates relative to the evaluation datasets using the same procedure as detailed in Section˜3.4, obtaining 9.3M images. We finetuned B-32 and B-16 CLIP models trained on DataComp-1B on our dataset to compare finetuning and pretraining performance. We also evaluate the original OpenAI CLIP [1], models pretrained on DataComp-M/L/1B, CommonPool-M/L [6], and DFN-5B [7], as well as the biological domain expert model BioCLIP [18], a ViT-B-16 CLIP model finetuned from OpenAI-CLIP on the TreeOfLife-10M dataset.
@@ 4.3 Object classification evaluation
To test the VLMs on object classification, we use the same procedure as CLIP [1], see the supplementary material for a detailed description. We evaluate all models on just encoding the class name, and on using the average embedding of the 80 context prompts that the CLIP authors used for ImageNet, and report the higher top-1 accuracy. For zero-shot object classification, we require models not to have been trained on the training set of the benchmark, to test “generalization to unseen datasets” [1]
@@
Benchmarks in the generic domain. We evaluate on ImageNet [14], a popular image classification benchmark [37]. We use the ILSVRC2012 validation set, which contains 50,000 images from 1,000 classes. The classes include simple objects, such as park bench, as well as more fine-grained labels like 23 types of terrier dogs, e.g., Staffordshire Bull Terrier. We further evaluate the robustness under distribution shifts on ImageNet-A [38], ImageNet-R [39], ImageNet-Sketch [40], ImageNet-V2 [41], and ObjectNet [42] as proposed by Taori et al. [43]. ImageNet-A contains 7500 samples of 200 ImageNet classes. The samples were adversarially filtered to make ResNet-50s misclassify them, providing a more challenging test. ImageNet-R contains 30000 renditions, such as paintings or embroidery, of 200 ImageNet object classes. ImageNet-Sketch contains 50000 sketches, covering 200 ImageNet classes. ImageNet-V2 replicates the original ImageNet generation process, providing an additional 10000 test images. The object-centered ObjectNet contains 18574 images from 113 ImageNet classes with control over background, rotation and viewpoint.
@@
Benchmarks in the expert domain. We evaluate our models on iNaturalist 2021 [20], a fine-grained species classification benchmark that contains 100k images in the validation set of 10k different species. Similar to Parashar et al. [44], we report the best results after testing on both the English common name and the Latin taxon name. We further test on the Caltech-UCSD Birds (CUB) [45] dataset, which contains 5,794 images of birds in the original author’s test set, each annotated as one of 200 fine-grained bird species, e.g., grasshopper sparrow. Additionally, we evaluate on the Rare Species benchmark proposed by Stevens et al. [18], comprising 400 species with 30 images each, specifically tailored to assess generalization to unseen taxa. To comply with the benchmark requirements of not seeing the testing 400 species during training, we exclude all entities and queries from our dataset that appear in RareSpecies, using substring matching. As class names, we evaluate all text types proposed by Stevens et al. [18]: combinations of the Latin taxonomy and the English common name. Same as in Section˜4.3 we evaluate on both the CLIP ImageNet prompt and no prompt, and report the better of both accuracies.